OpenStudy (anonymous):

Charlie and Juana are purchasing a townhouse and finance $94,700 with a 15-year 5/2 ARM at 7.35% with a 2/10 cap structure. What will their payments be at the beginning of the sixth year assuming they are charged the maximum interest rate for that year? $763.60 $677.53 $948.44 $869.83

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

@sydnoelle

ganeshie8 (ganeshie8):

u knw her ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes haha

ganeshie8 (ganeshie8):

okay :) both are asking similar questions lol

OpenStudy (anonymous):

I have a few other questions too!

OpenStudy (amistre64):

what does the given information tell us about refinancing times?

OpenStudy (anonymous):

should i put my question up here. or open a new one @ganeshie8

OpenStudy (amistre64):

in other words, what does the 5/2 ... 2/10 information tell us?

OpenStudy (anonymous):

That's where I'm confused. What are the ARMS's and cap structures.

OpenStudy (amistre64):

5/2 tells us about years .... 5 years at the intial rate, then it changes every 2 years

OpenStudy (amistre64):

2/10 tells us about interest changes ... it changes by 2% each time, not to exceed a total of 10%

OpenStudy (anonymous):

Oh okay

OpenStudy (amistre64):

so, the first thing I would do is determine the payment for a 15 year fixed rate loan, and balance it out after 5 years to determine the refinanced payments for the balance over the remaining 10 years

OpenStudy (anonymous):

How do I do that?

OpenStudy (amistre64):

how do you determine the payments of a 15 year fixed rate loan? this should be in your material

OpenStudy (amistre64):

I have my own formula that differs from the standard text formula

OpenStudy (amistre64):

ill show you mine if you show me yours lol

OpenStudy (anonymous):

\[M= P \times (i(1+i)^{nt} / (1+i) ^{nt} -1)\]

OpenStudy (anonymous):

Could you show my the answer? My computer is about to die. I think the answer would be C = $948.44

OpenStudy (anonymous):

By using that formula

OpenStudy (amistre64):

I use a basic structure:\[A=Bk^n+P\frac{1-k^n}{1-k}\] P is the payment we are looking for, B is the loan amount, A is what we oweand k and n are compounding factors; when A=0, the loan is paid off \[0=Bk^n+P\frac{1-k^n}{1-k}\]solve for P \[-Bk^n=P\frac{1-k^n}{1-k}\] \[-Bk^n\frac{1-k}{1-k^n}=P\] soo \[P=-94700(1+.0735/12)^{12*15}(1-(1+.0735/12))/(1-(1+.0735/12)^{12*15})\] which the wolf helps determine

OpenStudy (amistre64):

what payment amount for the first 5 years do you get?

OpenStudy (anonymous):

That isn't working in the calcualtor

OpenStudy (amistre64):

we can use this to find the amount owed after 5 years to refianance at and extra 2% interest for the remaining 10 years

OpenStudy (anonymous):

Oh! Okay so now we do the interest rate 2% more?

OpenStudy (amistre64):

lets find the balance after 5 years that needs to be refinanced first \[A=94700k^{12*5}-869.828\frac{1-k}{1-k^{12*5}};k=1+.0735/12\] http://www.wolframalpha.com/input/?i=94700k%5E%7B12*5%7D-869.828%5Cfrac%7B1-k%7D%7B1-k%5E%7B12*5%7D%7D%3Bk%3D1%2B.0735%2F12 im getting a balance of 136,593 to refinance for .0935 over 10 years. whats the payment for that type of fixed loan?

OpenStudy (anonymous):

$763.60?

OpenStudy (amistre64):

is that a guess, or did your formula produce that?

OpenStudy (anonymous):

I used the formula

OpenStudy (amistre64):

then that should be it :)

OpenStudy (anonymous):

Isn't your formula showing the same?

OpenStudy (amistre64):

maybe ... maybe not lol i got a typo in mine and its showing 1756 :)

OpenStudy (anonymous):

That is not even a choice

OpenStudy (anonymous):

@ganeshie8 can you help me like you helped Syd

OpenStudy (anonymous):

@amistre64 thats not an answer choice

OpenStudy (amistre64):

ARMS are not great for reducing payments, and the starting payments are 870, im thinking the reading of the phrase "beginning of the 6th year" might refer to the original payments of 870 instead of the adjusted payments to be made during the year

OpenStudy (amistre64):

ive dbl chked my thought process, its good. ARMs refinance the remaining balance for the remianing time period at a new interest rate which is why they are not the best option for home buyers.

OpenStudy (anonymous):

So what would the appropriate answer be?

OpenStudy (amistre64):

134212 is the remaining balance to play with .... had my fraction flipped

OpenStudy (amistre64):

im gonna wake up soon ... \[A=94700(1+\frac{.0735}{12})^{12*5}-869.83\frac{1-(1+\frac{.0375}{12})^{12*5}}{1-(1+\frac{.0375}{12})}=$73,763.90\] its these fat fingers on this tiny keyboard :)

OpenStudy (amistre64):

which is what my gut was saying since ARMs tend to have higher payments as they move on :)

OpenStudy (anonymous):

That was what I said in the beginning!

OpenStudy (amistre64):

you said the 700 one lol

OpenStudy (anonymous):

Scroll up to the very top

OpenStudy (amistre64):

all the way up there? thats just torture

ganeshie8 (ganeshie8):

lol this question should not be given in exam :/

OpenStudy (amistre64):

yes, thats the one you said to start with ;)